# On local Galois deformation rings: generalised tori

### v1

@@ 5. Admissible representations
This terminology is motivated by admissible Galois representations into $L$-groups appearing in the Langlands correspondence.
@@
(30) $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\xrightarrow{\cong}\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$
@@
The action of $\mathfrak{D}(M)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ factors through the action of $\mathfrak{D}(I_{\Delta}M)$, which acts freely.
@@
Proposition 5.2 implies that it is enough to prove the statement, when $A=\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$ and $\rho=\Phi_{\mathrm{nat}}\rho_{0}$.
@@ 6. Lafforgue’s $G$-pseudocharacters
We keep the notation of Section 5. We now recall Lafforgue’s notion of $G$-pseudocharacter in the form of [12, Definition 3.1]. The difference to Lafforgue’s original definition [8, Section 11] is that we work over the base ring $\mathcal{O}$ and allow $G$ to be disconnected. The definition works for generalised reductive $\mathcal{O}$-group schemes as defined in [11], but this generality is not needed here.
@@
We denote the set of $G$-pseudocharacters of $\Gamma_{1}$ over $A$ by $\mathrm{PC}_{G}^{\Gamma_{1}}(A)$. The functor $A\mapsto\mathrm{PC}_{G}^{\Gamma_{1}}(A)$ is representable by an affine $\mathcal{O}$-scheme $\mathrm{PC}_{G}^{\Gamma_{1}}$ [12, Theorem 3.20]. When $\varphi:G\to H$ is a homomorphism of generalised tori over $\mathcal{O}$, the induced maps $\varphi^{*}_{n}:\mathcal{O}[H^{n}]^{H^{0}}\to\mathcal{O}[G^{n}]^{G^{0}}$ give rise to an $H$-pseudocharacter $(\Theta_{n}\circ\varphi^{*}_{n})_{n\geq 1}$. By analogy with the notation for representations we denote this $H$-pseudocharacter by $\varphi\circ\Theta$. So we also have an induced map $\mathrm{PC}_{G}^{\Gamma}(A)\to\mathrm{PC}_{H}^{\Gamma}(A)$. It is easy to verify that specialisation along $f:A\to B$ commutes with composition with $\varphi$, i.e. $(\varphi\circ\Theta)\otimes_{A}B=\varphi\circ(\Theta\otimes_{A}B)$.
@@
We will show, that (37) and (39) are isomorphisms. The result is specific to generalised tori [^2] We don’t expect (37) and (39) to be isomorphisms when $G$ is an arbitrary generalised reductive group in the sense of [12, Definition 2.3]. However, we don’t know of an example where (37) is not an isomorphism. and might be of independent interest.
@@
As $G^{0}$ is linearly reductive the proof of [7, Proposition 2.11 (i)] applies, we recall the argument here. Let $\mathcal{O}(G^{\Gamma_{1}})$ be the $\mathcal{O}$-algebra which represents the functor
@@
We have a natural surjection $\mathcal{O}(G^{\Gamma_{1}})\twoheadrightarrow\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})$ with kernel $J$. We also have a natural surjection $\mathcal{O}(G^{\Gamma_{1}})^{G^{0}}\twoheadrightarrow\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})$ with kernel $I$. The structure of $J$ and $I$ is described explicitly in [7, Proposition 2.5]; note that there $G$ is assumed to be connected, but the description generalises easily to our situation. Namely $J$ is generated by the image of the $G^{0}$-equivariant maps
@@
by 6.3. We conclude, that (41) is obtained from (40) by tensoring with $-\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$. So by 6.2 (41) is an isomorphism as well. ∎
@@ 7.1. Deformations of representations
Conversely, if $I$ is an ideal of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})_{\mathfrak{m}}$ such that the quotient $A$ is finite then $A\in\mathfrak{A}_{\mathcal{O}}$. By specialising the universal representation $\Gamma_{1}\rightarrow G(\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}))$ along $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow A$ we obtain a representation $\rho:\Gamma_{1}\rightarrow G(A)$. Since the target is finite $\rho$ extends uniquely to a continuous representation $\tilde{\rho}:\widehat{\Gamma}_{1}\rightarrow G(A)$, which defines a point in $D^{\square}_{\overline{\rho}}(A)$.
@@
If $N$ is finitely generated as $\mathbb{Z}_{p}$-module, then $N\cong\mu\oplus\mathbb{Z}_{p}^{r}$, where $\mu$ is a finite $p$-group and $r=\dim_{\mathbb{Q}_{p}}N\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$. Hence, $\mathcal{O}\llbracket N\rrbracket\cong\mathcal{O}[\mu]\llbracket x_{1},\ldots,x_{r}\rrbracket$ and hence (1) implies (3).
@@
Proposition 5.2 allows us to identify $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$. It follows from (17) that if $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is finite then $\mathcal{E}/p\mathcal{E}$ is also finite. Since $M$ is a finite free $\mathbb{Z}$-module, the assertion follows from Lemma 7.3 applied to $\mathcal{A}=(\mathcal{E}\otimes M)_{\Delta}$. ∎
@@
(46) $h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})\leq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}-\dim G_{k}$
@@
implies that $r=h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})+\dim G_{k}$. Let $N:=((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$. By considering (47) modulo $\varpi$ and using Lemma 7.5 we deduce that $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N\leq r$. Moreover, the assumption $r-s\leq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ implies that $r-s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ and $\varpi,f_{1},\ldots,f_{s}$ can be extended to a system of parameters of a regular ring $\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket$. Thus $\varpi,f_{1},\ldots,f_{s}$ are a part of a regular sequence, and hence $R^{\square}_{\overline{\rho}}$ is complete intersection, flat over $\mathcal{O}$ of relative dimension of $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$. ∎
@@ 7.2. Deformations of pseudocharacters
As the image of $\Gamma_{1}$ in $\widehat{\Gamma}_{1}$ is dense, injectivity follows from [12, Lemma 3.2]. Let $\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$. By 6.6 there is a representation $\rho\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$, such that $\Theta=\Theta_{\rho}$. Since $A$ is finite, $\rho$ extends to a continuous representation $\tilde{\rho}:\widehat{\Gamma}_{1}\to G(A)$ and we have $\Theta_{\tilde{\rho}}|_{\Gamma_{1}}=\Theta$. ∎
Let $\overline{\Theta}$ be the $G$-pseudocharacter of $\widehat{\Gamma}_{1}$ associated to $\overline{\rho}$. Let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the deformation functor which sends $A$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\widehat{\Gamma}_{1}$ valued in $A$ such that $\Theta\otimes_{A}k=\overline{\Theta}.$ The functor $D^{\mathrm{ps}}_{\overline{\Theta}}$ is pro-represented by $R^{\mathrm{ps}}_{\overline{\Theta}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ by [12, Theorem 5.4].
@@
Since $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty Corollary 6.5 allows us to identify $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[H_{1}(\Gamma,M)]$. The assertion follows from Proposition 7.9 and Lemma 7.2. ∎
@@
If $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is a finite $p$-group and $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian and is naturally isomorphic to the $\mathfrak{m}$-adic completion of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$, where $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ corresponding to $\overline{\Theta}|_{\Gamma_{1}}$.
@@ 7.4. Irreducible components
This is a special case of Lemma 7.2 applied with $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$ and $\overline{\rho}:\Gamma_{F}\rightarrow G(k),\gamma\mapsto(1,\pi(\gamma))$ via Proposition 5.2 with $\rho_{0}:\Gamma_{F}\rightarrow G(\mathcal{O}),\gamma\mapsto(1,\pi(\gamma))$. ∎
@@
Since $A^{\mathrm{gen}}_{G,\overline{\Theta}}\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$ by 7.13 the map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is $\mathrm{X}(\mu)$-equivariant and induces a an $\mathrm{X}(\mu)$-equivariant bijection between the sets of irreducible components.

### v2

@@ 5. Admissible representations
This terminology is motivated by admissible Galois representations into $L$-groups appearing in the Langlands correspondence, see [3, §9].
@@
(30) $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\xrightarrow{\cong}\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}].$
@@
Assume that there is a representation $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$. The action of $\mathfrak{D}(M)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ factors through the action of $\mathfrak{D}(I_{\Delta}M)$, which acts freely.
@@
Let $A$ be any $\mathcal{O}$-algebra and let $H(A)$ be the group $\operatorname{Hom}((\mathcal{E}\otimes M)_{\Delta},A^{\times})$. Since $q:(\mathcal{E}\otimes M)_{\Delta}\rightarrow I_{\Delta}M$ is surjective, we may identify $\mathfrak{D}(I_{\Delta}M)(A)$ with a subgroup of $H(A)$. The isomorphism (30) identifies $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ with $H(A)$ and the action of $\mathfrak{D}(I_{\Delta}M)(A)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ is identified with the action of $\mathfrak{D}(I_{\Delta}M)(A)$ on $H(A)$ by multiplication, which is a free action. Hence, the action of $\mathfrak{D}(I_{\Delta}M)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ is free. ∎
@@
Proposition 5.2 implies that it is enough to prove the statement, when $A=\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$ and $\rho=\Phi_{\mathrm{nat}}\rho_{0}$, where $\rho_{0}$ is any representation in $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ and $\Phi_{\mathrm{nat}}$ is the cocycle defined in Section 4.
@@ 6. Lafforgue’s $G$-pseudocharacters
We keep the notation of Section 5. We now recall Lafforgue’s notion of $G$-pseudocharacter in the form of [13, Definition 3.1]. The difference to Lafforgue’s original definition [9, Section 11] is that we work over the base ring $\mathcal{O}$ and allow $G$ to be disconnected. The definition works for generalised reductive $\mathcal{O}$-group schemes as defined in [12], but this generality is not needed here.
@@
We denote the set of $G$-pseudocharacters of $\Gamma_{1}$ over $A$ by $\mathrm{PC}_{G}^{\Gamma_{1}}(A)$. The functor $A\mapsto\mathrm{PC}_{G}^{\Gamma_{1}}(A)$ is representable by an affine $\mathcal{O}$-scheme $\mathrm{PC}_{G}^{\Gamma_{1}}$ [13, Theorem 3.20]. When $\varphi:G\to H$ is a homomorphism of generalised tori over $\mathcal{O}$, the induced maps $\varphi^{*}_{n}:\mathcal{O}[H^{n}]^{H^{0}}\to\mathcal{O}[G^{n}]^{G^{0}}$ give rise to an $H$-pseudocharacter $(\Theta_{n}\circ\varphi^{*}_{n})_{n\geq 1}$. By analogy with the notation for representations we denote this $H$-pseudocharacter by $\varphi\circ\Theta$. Thus we also have an induced map $\mathrm{PC}_{G}^{\Gamma}(A)\to\mathrm{PC}_{H}^{\Gamma}(A)$. It is easy to verify that specialisation along $f:A\to B$ commutes with composition with $\varphi$, i.e. $(\varphi\circ\Theta)\otimes_{A}B=\varphi\circ(\Theta\otimes_{A}B)$.
@@
We will show, that (37) and (39) are isomorphisms. The result is specific to generalised tori [^2] We don’t expect (37) and (39) to be isomorphisms when $G$ is an arbitrary generalised reductive group in the sense of [13, Definition 2.3]. However, we don’t know of an example where (37) is not an isomorphism. and might be of independent interest.
@@
As $G^{0}$ is linearly reductive the proof of [8, Proposition 2.11 (i)] applies, we recall the argument here. Let $\mathcal{O}(G^{\Gamma_{1}})$ be the $\mathcal{O}$-algebra which represents the functor
@@
We have a natural surjection $\mathcal{O}(G^{\Gamma_{1}})\twoheadrightarrow\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})$ with kernel $J$. We also have a natural surjection $\mathcal{O}(G^{\Gamma_{1}})^{G^{0}}\twoheadrightarrow\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})$ with kernel $I$. The structure of $J$ and $I$ is described explicitly in [8, Proposition 2.5]; note that there $G$ is assumed to be connected, but the description generalises easily to our situation. Namely $J$ is generated by the image of the $G^{0}$-equivariant maps
@@
by 6.3. We conclude, that (41) is obtained from (40) by applying $-\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$. So by 6.2 (41) is an isomorphism as well. ∎
@@ 7.1. Deformations of representations
Conversely, if $I$ is an ideal of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})_{\mathfrak{m}}$ such that the quotient $A$ is finite then $A\in\mathfrak{A}_{\mathcal{O}}$. The map $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow k$ corresponding to $\overline{\rho}$ factors through $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\rightarrow k$ as $\overline{\rho}|_{\Gamma_{1}}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(k)$ by assumption. Since $G/G^{0}$ is constant and $A$ is a local ring we have $(G/G^{0})(A)=(G/G^{0})(k)$ and hence the map $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow A$ factors through $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\rightarrow A$. By specialising the universal representation $\Gamma_{1}\rightarrow G(\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}))$ along this map we obtain a representation $\rho:\Gamma_{1}\rightarrow G(A)$. Since the target is finite $\rho$ extends uniquely to a continuous representation $\tilde{\rho}:\widehat{\Gamma}_{1}\rightarrow G(A)$, which defines a point in $D^{\square}_{\overline{\rho}}(A)$.
@@
If $N$ is finitely generated as a $\mathbb{Z}_{p}$-module, then $N\cong\mu\oplus\mathbb{Z}_{p}^{r}$, where $\mu$ is a finite $p$-group and $r=\dim_{\mathbb{Q}_{p}}N\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$. Hence, $\mathcal{O}\llbracket N\rrbracket\cong\mathcal{O}[\mu]\llbracket x_{1},\ldots,x_{r}\rrbracket$ and hence (1) implies (3).
@@
Proposition 5.2 allows us to identify $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$. It follows from (17) that if $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is finite then $\mathcal{E}/p\mathcal{E}$ is also finite. Since $M$ is a finite free $\mathbb{Z}$-module, the assertion follows from 7.1 and Lemma 7.3 applied to $\mathcal{A}=(\mathcal{E}\otimes M)_{\Delta}$. ∎
@@
(46) $h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})\geq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}-\dim G_{k}$
@@
implies that $r=h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})+\dim G_{k}$. Let $N:=((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$. By considering (47) modulo $\varpi$ and using Lemma 7.5 we deduce that $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N\geq r-s$. Moreover, the assumption (46) implies that $r-s\geq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$. Hence, $r-s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ and $\varpi,f_{1},\ldots,f_{s}$ can be extended to a system of parameters of a regular ring $\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket$. Thus $\varpi,f_{1},\ldots,f_{s}$ are a part of a regular sequence, and hence $R^{\square}_{\overline{\rho}}$ is complete intersection, flat over $\mathcal{O}$ of relative dimension of $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$. ∎
@@ 7.2. Deformations of pseudocharacters
As the image of $\Gamma_{1}$ in $\widehat{\Gamma}_{1}$ is dense, injectivity follows from [13, Lemma 3.2]. Let $\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$. By 6.6 there is a representation $\rho\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$, such that $\Theta=\Theta_{\rho}$. Since $A$ is finite, $\rho$ extends to a continuous representation $\tilde{\rho}:\widehat{\Gamma}_{1}\to G(A)$ and we have $\Theta_{\tilde{\rho}}|_{\Gamma_{1}}=\Theta$. ∎
Let $\overline{\Theta}$ be the $G$-pseudocharacter of $\widehat{\Gamma}_{1}$ associated to $\overline{\rho}$. Let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the deformation functor which sends $A$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\widehat{\Gamma}_{1}$ valued in $A$ such that $\Theta\otimes_{A}k=\overline{\Theta}.$ The functor $D^{\mathrm{ps}}_{\overline{\Theta}}$ is pro-represented by $R^{\mathrm{ps}}_{\overline{\Theta}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ by [13, Theorem 5.4].
@@
Since $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty Corollary 6.5 allows us to identify $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[H_{1}(\Gamma,M)]$. The assertion follows from Proposition 7.9 and Lemma 7.2. ∎
@@
If $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is a finite $p$-group and $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian and is naturally isomorphic to the $\mathfrak{m}$-adic completion of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$, where $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ corresponding to $\overline{\Theta}|_{\Gamma_{1}}$.
@@ 7.4. Irreducible components
Let $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$, $\overline{\rho}:\widehat{\Gamma}_{1}\rightarrow G(k),\gamma\mapsto(1,\pi(\gamma))$ and let $\rho_{0}:\widehat{\Gamma}_{1}\rightarrow G(\mathcal{O}),\gamma\mapsto(1,\pi(\gamma))$. The map $\Phi\mapsto\Phi\rho_{0}$ induces a natural bijection between $\widehat{Z}^{1}(A)$ and $D^{\square}_{\overline{\rho}}(A)$ for all $A\in\mathfrak{A}_{\mathcal{O}}$; this can be shown by the same argument as in the proof of 5.2. Thus $\widehat{Z}^{1}$ is pro-represented by $R^{\square}_{\overline{\rho}}$ and the assertion follows from 5.2, 7.1 and 7.2. ∎
@@
Since $A^{\mathrm{gen}}_{G,\overline{\Theta}}\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$ by 7.13 the map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is $\mathrm{X}(\mu)$-equivariant and induces an $\mathrm{X}(\mu)$-equivariant bijection between the sets of irreducible components.